# Quantum Sensing Using Cold Atoms and Molecules for Nuclear and Particle Physics

### v2

@@ VIII.2 Experimental results
Since the publications of theory papers around 2018, a large number of experimental results has been reported on this topic, ranging from isotope shift measurements alone to direct constraints on the existence of new forces. For certain atomic species, such as Xe [254], Cd [358], and Nd [359], isotope shifts were reported in the context of the search for new bosons. The constraints on the existence of new bosons are summarized in Fig. 9. The first experimental constraints on the new bosons were established through isotope shift measurements for Yb+ [348] and Ca+ [344]. In these studies, isotope shift measurements of Ca+ with a 20-Hz uncertainty excluded the existence of new bosons based on the observed linearity of the King plot, whereas nonlinearity is observed in the isotope shift measurements of Yb+ with a 300-Hz uncertainty. The initial Yb+ study focused on two E2 transitions, and later measurements included isotope shifts for the E3 transition with an accuracy of 500 Hz[346]. Additionally, multiple transitions in neutral Yb were carefully analyzed [347, 349, 122]. The most recent results for the E2 and E3 transitions of Yb+ achieved uncertainties of 5 and 16 Hz, respectively [350]. While Ca+ and some Yb measurements did not observe nonlinearity, consistent with constraint based on the measurement for electrons and neutron scattering, certain Yb dataset exhibit nonlinearity , leading to allowed regions for the new bosons. These allowed and excluded regions contradict each other, suggesting that the observed nonlinearity may be induced by some background effects within the SM. This calls for further investigation. A recent observation of nonlinearity in Ca [255] was based on the combination of the precision spectroscopy of Ca+ and Ca14+. This study attributed the observed nonlinearity to the second order mass shift and nuclear polarization, setting most stringent constraint on the new bosons.
@@
When nonlinearity is observed in the King plot, its source must be carefully analyzed. For this purpose, nonlinearity is decomposed into the following two components: , where is the residual of four data points from a linear fit [348] (Note that in Ref. [346] is equivalent to ). On the plane, a single source of nonlinearity with varying strength forms a line passing through the origin. When residuals of King plots for different combination of transitions in Yb are plotted on this plane, the data points approximately align along a single line. This suggests that the observed nonlinearity primarily originates from a single source, which can be explained by nuclear density functional theory calculations. These calculations incorporate several factors, including the surface thickness of nuclear density, relativistic corrections, nuclear deformation, and pairing effects [346]. The isotope shift measurements are so precise that they can select a nuclear model that accurately describes the experimental results. [346, 350].
@@
IX Atomic parity violation
IX.1 Theoretical basis
Parity (P) transformation refers to flipping the sign of one spatial coordinate. In three dimensions, this is equivalent to the transformation . Under this transformation, fundamental equations of motion, such as the Newton’s equation of motion, the Schrödinger equation, and the Dirac equation, remain unchanged. This invariance under P transformation is known as P symmetry. However, among the four fundamental forces, P symmetry is completely violated in the weak force. Historically, possibility of P violation was proposed by T.D. Lee and C. N. Yang [363]. Shortly thereafter, C. S. Wu experimentally confirmed it [364]. The discovery was later incorporated into the SM as a part of the weak interaction theory.
In atoms, where the electromagnetic force dominates, P symmetry is naively conserved. If the P symmetry is violated in interactions in atoms, it is induced by the weak force or BSM physics. Examples of such interactions are summarized in the Feynman diagrams in Fig. 10. Fig.10(a) shows an interaction between an electron and a nucleus mediated by a Z boson instead of a photon. Although this interaction occurs at tree-level, its effect is very small due to the large mass of the Z boson. Fig.10(b) and (c) depict higher-order perturbative processes, where an additional Z boson mediating the force, or a virtual W or Z boson forms a loop within the nucleus. If a hypothetical boson appears in a loop or mediates the force between the nucleus and the electron, this interaction would indicate the presence of BSM physics[365, 366]. The effective Hamiltonian for spin-independent APV is given by
where is the Fermi constant for the weak force, with being the Dirac matrices, and () is the creation (annihilation) operator for an electron. The effective weak charge consists of the SM contribution and the BSM contribution .
with and being number of protons and neutrons, respectively, being the Weinberg angle, and being the BSM contribution to the vector current, where are the coupling of the BSM physics to up and down quarks, respectively. This equation indicates that APV experiments effectively measure . The energy scale of this interaction corresponds to the size of the hydrogen-like 1S orbital for a nucleus with charge , i.e., the innermost electronic orbital. This leads to a momentum transfer of [365]. Intuitively, this occurs because electrons experience the weak force primarily in regions where the Coulomb potential from the nucleus is unscreened. This momentum transfer is significantly lower than that in other measurements of performed in scattering experiments. Searches for deviations of the from serves as a probe for P-violating BSM physics, e.g., Z’ boson and leptoquarks [365, 367, 368].
Spin-dependent APV originates from all three Feynman diagrams shown in Fig. 10. Diagram (c) illustrates the contribution of the nuclear anapole moment, which is analogous to a solenoidal current bent into a toroidal shape. The anapole moment is proportional to nuclear spin , and serves as a probe for nuclear parity violation. Some constraints on the anapole moment were set through the Cs experiment discussed in the following section.
The effect of P violation can be observed as a nonzero transition amplitude for an electronic transition between states of the same parity. P violation causes mixing between states of opposite parity, which slightly allows transitions between the same parity states, something normally forbidden in E1 transitions. In practice, measuring APV involves evaluating the contribution of the P-violating term in the total matrix element of the otherwise forbidden transition between electronic orbitals of the same parity.
Because the effect of APV is extremely small, various enhancement techniques are used to amplify the signal to an observable level. The mixing parameter between an state and a state due to APV satisfies the following equation.
where is the transition energy for the transition of interest. Due to the scaling [369, 370], heavy atoms are preferred for APV measurements. The measured quantity is the interference term between the P-conserving amplitude and the P-violating amplitude :
where is the dc electric field applied to mix parity-odd and parity-even states, is the tensor transition polarizability, is the handedness of the polarization of the oscillating field, is the magnitude of the oscillating field perpendicular to , and are Clebsch-Gordan coefficients, and is the E1 transition matrix element due to APV [371]. is extracted from experimental measurements and converted to and .
IX.2 Experimental reports
The two-level system suitable for APV consists of the ground state and an excited state with the same parity as the ground state. Due to the enhancement, heavy atoms are preferred, and historically measurements were mainly performed using Cs atoms, as shown in Table 7. In these experiments, the 6S ground state and the 7S excited state are regarded as the two-level system. The E1 transition between these two state is forbidden, but they can be coupled via an M1 transition. In the experiment, a standing wave generated by a build-up cavity helps suppress the M1 transition. To drive the 6S-7S transition in a controlled manner, the 7P state is mixed with the 7S state by a static electric field ( in Eq. IX.1). By flipping the sign of and , additional discrimination between the backgrounds and the signal is achieved. After accounting for all systematic uncertainties, most of which are related to the M1 amplitude , is obtained[372], which leads to around 2.4 MeV.
Another attempt to observe APV is performed using Yb. In this case, the two-level system consists of the ground state and the state. A significant parity-violating effect is observed [373] for this transition. One advantage of Yb over Cs is the number of stable isotopes. While Cs has only one, Yb has five spinless isotopes. This allows for the estimation of neutron contribution to by measuring the P-violating effect across multiple isotopes [374].
The next generation experiment aims to measure the APV in Fr. Because all Fr isotopes are unstable, these experiments are conducted at accelerator facilities. Fr atoms generated by nuclear reactions were successfully loaded into a magneto-optical trap in the 1990s at Stony Brook [375]. Since then, relevant transitions have been characterized in preparation for APV measurement[376, 377]. An effort to improve the accuracy of APV measurements is Cs is also ongoing [378].
Another method to enhance sensitivity is to use a small . To achieve this, the use of diatomic molecules [379] and dysprosium [380] is proposed. In these species, closely spaced states with the opposite parity enhances the mixing shown in Eq. IX.1. Experimental sensitivity has been reported for a BaF molecule experiment [381]. In the case of BaF, the measured quantity is the nuclear spin dependent APV. A MOT for BaF is recently realized [382].
X Electric dipole moment search
X.1 Theoretical basis
Electric dipole moments (EDMs) of elementary particles serve as a probe for high-energy physics through the violation of time-reversal (T) symmetry. If a spin-1/2 elementary particle has a nonzero EDM, it violates T symmetry because the EDM remains unchanged while the magnetic dipole moment flips under the T operation. Assuming the CPT theorem [386, 387] (C stands for the charge conjugation), this implies a violation of the symmetry under the CP operation. Such a CP-violating moment is predicted in the SM of particle physics. For the electron, four-loop perturbations predict an EDM of approximately cm [388]. denotes the EDM of , where is for electron or some nuclei such as Yb, Hg, and Ra. CP-violating interactions between electrons and nucleus mimics an electron EDM. This effect is estimated to range from cm [389] to cm [388]. In supersymmetric theories, CP-violating moments can arise from one-loop diagrams, resulting in significantly larger EDM values. This argument extends to other elementary particles and nuclei, making EDM searches an important avenue to test supersymmetric theories. Also, CP-violation is closely linked to the matter-antimatter asymmetry in the universe. Moreover, if an axion-like particle contributes via loop diagrams at a vertex, EDM searches can put constraints on such particles [390, 391].
The search for EDM began with a theoretical proposal in 1950 [392]. Since then, experiments have been conducted for various particles, including the neutron, electron, muon, and certain nuclei, as summarized in a separate review paper[393]. Here, since AMO physics is focused on, only searches for electron EDM and some nuclear EDM are discussed.
Schiff’s theorem [394] plays an important role in EDM searches. It states that the overall EDM of an atom vanishes if it consists of point-like nonrelativistic charged particles interacting with a static electric field. However, a measurable EDM can arise due to relativistic effects, the finite size of the nucleus, and the presence of an unpaired electron. As a result, searches for the electron EDM are always performed using atoms or molecules with an unpaired electron. The significance of relativistic effects and enhancement of the electron EDM in an atom, as discussed in Ref. [395], forms the foundation for EDM searches using atoms and molecules.
In experiments for EDM searches, the small effect of the EDM is investigated using a Ramsey sequence under both a magnetic field and an electric field . If a particle only has a magnetic moment due to its spin, it precesses according to the Hamiltonian . If the particle possesses a nonzero EDM , the amount of precession changes according to the modified Hamiltonian .
X.2 Electron EDM
The electron is one of the elementary particles for which EDM searches have long been performed, as shown in Fig. 11. Because it is a lepton, BSM physics contributions at single-loop level arises from new particles with leptonic quantum number, e.g., the supersymmetric partner of the electron [396]. The electron EDM was first investigated indirectly by examining its effect on the Lamb shift in hydrogen atoms [397]. Although the first direct search was performed with bare electrons [398], modern experiments primarily use neutral atoms and molecules. This approach allows for the application of stronger effective electric fields, enhancing the EDM-induced precession through internal electric fields.
For an atom with atomic number , the enhancement factor of the electron EDM is [423]. Because of this dependence, heavy atoms are typically used in EDM searches. Most of the measurements with atoms shown in Fig. 11 were performed with Cs and Tl.
To achieve even stronger internal electric fields, polar molecules are utilized. Molecules composed of a soft, heavy atom and a highly electronegative atom can generate an internal electric field on the order of 10 GV/cm [424, 425, 426]. Molecules are selected such that they have one additional electron beyond fully closed inner shells and valence electrons involved in chemical bonding. This extra electron serves as the probe for EDM measurements. The YbF experiment [416] pioneered the use of molecules in EDM searches, and recent progresses are made with ThO [421] and HfF+ molecules [422].
In these molecules, a set of states known as -doublet states is used for EDM searches. is the projection of the electron’s total angular momentum on the internuclear axis of the diatomic molecule. For the case of , states are nearly degenerated, forming two closely spaced states with parity and [427]. The -doublet can be fully polarized by applying a small external electric field. When polarized, a pair of states at the same energy exhibits the same orientation but opposite spin orientation. The Ramsey sequence is performed under external magnetic and electric fields. The measured signal corresponds to the precession of the state relative to the state. To cancel systematic shifts and extract the EDM signal, the experiment involves flipping external electric and magnetic fields, as well as the orientation of the molecule. The difference in precession across all combinations of these flips are systematically processed to extract the desired EDM signal.
The figure of merit for the sensitivity follows the same SQL as atomic clocks [428]:
where is the number of molecules including detection efficiency. The choice of molecule is based on and the coherence time of the -doublet states, which limits . For ThO, the coherence time is 4.2(5) ms [429]. Given this timescale, a beam experiment is suitable, and measurements are conducted using buffer gas-cooled cryogenic beam[430]. It is relatively easy to increase the number of molecules to enhance the sensitivity by reduced QPN. In contrast, the -doublet states in HfF+ ion have significantly longer lifetime, allowing an interrogation time of up to 3 s [422]. Although the number of trapped ions is much smaller compared to molecules in a beam, this extended interrogation time compensates for this limitation. As a result, the sensitivity of these two state-of-the-art experiments are different by only a small factor.
Once a sufficient level of sensitivity is achieved, all systematic shifts must be accurately estimated and compensated to extract the electron EDM with small uncertainty. Two major sources of uncertainty are nonuniform magnetic fields and non-inverting components of the applied electric and magnetic fields. Another source of uncertainty arises from laser power stability, including variations in the overlap between the light and the molecules. Additionally, in the HfF+ experiment, where ions undergo circular motion, uncertainties specific to this motion also needs to be carefully considered.
The most stringent constraint on the electron EDM comes from the HfF+ experiment, which sets an upper limit of cm [422]. The ThO experiment provides the constraint of cm [421]. When properly interpreted within the framework of supersymmetric theories, these constraints correspond to TeV-scale physics[396]. Previous experiments have already excluded the naive supersymmetric theory, and the latest results constrain alternative theoretical frameworks. The highest attainable energy scale accessible through these constraints is 40 TeV [422].
Upgrades to the ThO experiment [431] and investigation into a successor experiment for the HfF+ experiment[432], as well as for improvements for the YbF experiment [433] are currently underway. At the proposal stage, enhancing sensitivity using entangled states is discussed. One such proposal suggests integrating a cavity QED system with an electron EDM measurement using Fr to generate a squeezed-spin state for Fr [434]. Experimental efforts are also focused on future EDM searches with radioactive isotopes, particularly in molecules. RaF has gained a significant interest not only for electron EDM searches but also for APV studies[435]. Experimental investigations of molecular spectroscopy of RaF are ongoing [436]. In general, Ra-containing molecules attract considerable attention, with both theoretical[437, 438] and experimental[439] studies being conducted on various species. In addition, polyatomic molecules are being explored as promising next-generation candidates for electron EDM searches. A theoretical proposal [440] pointed out that certain polyatomic molecules have states easy to polarize, analogous to doublets, yet independent of electronic spins. This property prevents spin-related complications in laser cooling and sensitivity to the electron EDM. This proposal triggered investigations on YbOH [441] and YbOCH3 [442]. Recently, a MOT for SrOH was demonstrated [443] in the context of electron EDM searches. Furthermore, a theoretical proposal suggests that the sensitivity of EDM searches with polyatomic molecules could be enhanced through quantum entanglement [444].
X.3 Nuclear EDM
Within the technology of AMO physics, nuclear EDMs can also be measured, with current constraints reaching the 10-TeV scale physics. Nuclear EDM measurements are performed using atomic or molecular systems where electrons are fully paired to eliminate any effects from the electron EDM. Nuclei with spin 1/2 are preferred due to the simple structure and the absence of deformation effects. As a result, popular species for nuclear EDM searches are 129Xe, 171Yb, 199Hg, 225Ra, and the neutron. To measure a nuclear EDM, spin-polarized nuclei are used. One method to prepare and detect such a spin-polarized system is through optical pumping and optical detection with lasers. Among the species listed above, experiments with 171Yb, 199Hg, and 225Ra rely on such atomic physics techniques and are the focus here. For other systems, a detailed summary can be found in another review paper [393].
In the measurement sequence, nuclei are placed under magnetic and electric fields, and a Ramsey sequence is performed to detect the tiny phase difference when the direction of the electric field is reversed. The relevant energy level corresponds to the nuclear spin 1/2 in a strong magnetic field, meaning the energy scale for the Ramsey sequence is in the range of NMR spectroscopy, not laser spectroscopy. Because the internal electric field of the nucleus is zero, the applied electric field originates from electrodes, and is significantly smaller than in electron EDM searches. However, due to the exceptionally long coherence time of nuclear spin, the interrogation time can reach s, resulting in constraints of cm for 199Hg [445], cm [446] for 171Yb, and cm for 226Ra [447], which are comparable to the constraint on the electron EDM.
The Hg experiment uses a vapor cell, making the major systematic uncertainty related to atomic motion. Yb atoms are laser-cooled and trapped in an optical dipole trap, eliminating motion-related uncertainty. However, the optical dipole trap introduces other sources of uncertainty. The 225Ra experiment [447] has a similar configuration to the Yb experiment, with its source of uncertainty arising from a phase shift proportional to the square of the electric field and the magnetic field correlated to the electric field.
XI Summary and Outlook
@@
High precision measurements with quantum states in atoms can detect small energy shifts induced by phenomena originating from nuclear and particle physics. For future advancements, several key aspects are important. One is improvement in measurement precision. Even within the same experimental system, increased measurement precision enhances sensitivity to nuclear and particle physics phenomena in most of the topics covered in this paper. In the past, improvements in clock uncertainty contributed to searches for ultralight dark matter and time variation of fundamental constants. Secondly, the use of highly sensitive quantum sensors can significantly improve the sensitivity. A notable example is searches for variation of the fine structure constant. Switching from conventional ion clocks to HCI clocks or nuclear clocks would enhance the sensitivity by orders of magnitude. In addition, a solid theoretical understanding of the measurements is crucial. Theoretical calculations of atomic or molecular structures determine the sensitivity of a quantum sensor to the quantity of interest. Examples include the sensitivity coefficient for the variation of the fine structure constant and the internal electric field for the electron EDM search. Further improvements in the accuracy of these theoretical calculations can refine the determination of parameters in nuclear and particle physics. Moreover, new theoretical proposals can inspire extensive experimental searches for previously unexplored parameter spaces. Historically, searches for the fifth force using isotope shifts and ultralight dark matter expanded following initial theoretical proposals. With the combination of these factors, further exploration of nuclear and particle physics using atoms and molecules as quantum sensors are expected.
@@ Acknowledgements.
- Cong et al. [2024] L. Cong, W. Ji, P. Fadeev, F. Ficek, M. Jiang, V. V. Flambaum, H. Guan, D. F. J. Kimball, M. G. Kozlov, Y. V. Stadnik, et al., “Spin-dependent exotic interactions,” (2024), arXiv:2408.15691 [hep-ph] .
@@
- McGrew et al. [2018a] W. F. McGrew, X. Zhang, R. J. Fasano, S. A. Schäffer, K. Beloy, D. Nicolodi, R. C. Brown, N. Hinkley, G. Milani, M. Schioppo, et al., “Atomic clock performance enabling geodesy below the centimetre level,” Nature 564, 87–90 (2018a).
@@
- McGrew et al. [2018b] W. F. McGrew, X. Zhang, R. J. Fasano, S. A. Schäffer, K. Beloy, D. Nicolodi, R. C. Brown, N. Hinkley, G. Milani, M. Schioppo, et al., “Atomic clock performance enabling geodesy below the centimetre level,” Nature 564, 87–90 (2018b).
@@
- Zheng et al. [2022a] X. Zheng, J. Dolde, V. Lochab, B. N. Merriman, H. Li, and S. Kolkowitz, “Differential clock comparisons with a multiplexed optical lattice clock,” Nature 602, 425–430 (2022a).
@@
- Jiang, Arora, and Safronova [2008] D. Jiang, B. Arora, and M. S. Safronova, “Electric quadrupole moments of metastable states of , , and ,” Phys. Rev. A 78, 022514 (2008).
@@
- Beeks et al. [2024] K. Beeks, G. A. Kazakov, F. Schaden, I. Morawetz, L. T. de Col, T. Riebner, M. Bartokos, T. Sikorsky, T. Schumm, C. Zhang, et al., “Fine-structure constant sensitivity of the Th-229 nuclear clock transition,” (2024), arXiv:2407.17300 [nucl-th] .
@@
- Geller [1970] R. Geller, “New high intensity ion source with very low extraction voltage,” Appl. Phys. Lett. 16, 401–404 (1970), https://pubs.aip.org/aip/apl/article-pdf/16/10/401/18423299/401_1_online.pdf .
@@
- Wilzewski et al. [2024] A. Wilzewski, L. I. Huber, M. Door, J. Richter, A. Mariotti, L. J. Spieß, M. Wehrheim, S. Chen, S. A. King, P. Micke, et al., “Nonlinear calcium king plot constrains new bosons and nuclear properties,” (2024), arXiv:2412.10277 [physics.atom-ph] .
@@
- Schaden et al. [2024] F. Schaden, T. Riebner, I. Morawetz, L. T. D. Col, G. A. Kazakov, K. Beeks, T. Sikorsky, T. Schumm, K. Zhang, V. Lal, et al., “Laser-induced quenching of the Th-229 nuclear clock isomer in calcium fluoride,” (2024), arXiv:2412.12339 [physics.atom-ph] .

### v3

@@ VIII.2 Experimental results
Since the publications of theory papers around 2018, a large number of experimental results has been reported on this topic, ranging from isotope shift measurements alone to direct constraints on the existence of new forces. For certain atomic species, such as Xe [270], Cd [381], and Nd [382], isotope shifts were reported in the context of the search for new bosons. The constraints on the existence of new bosons are summarized in Fig. 9. The first experimental constraints on the new bosons were established through isotope shift measurements for Yb+ [370] and Ca+ [366]. In these studies, isotope shift measurements of Ca+ with a 20-Hz uncertainty excluded the existence of new bosons based on the observed linearity of the King plot, whereas nonlinearity is observed in the isotope shift measurements of Yb+ with a 300-Hz uncertainty. The initial Yb+ study focused on two E2 transitions, and later measurements included isotope shifts for the E3 transition with an accuracy of 500 Hz[368]. Additionally, multiple transitions in neutral Yb were carefully analyzed. At first, isotope shifts for the SD2 transition is reported with kHz accuracy [371] . Then the isotope shifts for the conveitional SP0 clock transition are reported [369]. Because of its high accuracy on the order of 1 Hz, this data are often used for analyses in other reports. The S transition is first incorporated into a King-plot analysis in Ref. [125], with a recent improved report of accuracy [373]. The most recent results for the E2 and E3 transitions of Yb+ achieved uncertainties of 5 and 16 Hz, respectively [372]. While Ca+ and some Yb measurements did not observe nonlinearity, consistent with the constraint based on the measurement for electrons and neutron scattering, certain Yb dataset exhibit nonlinearity, leading to allowed regions for the new bosons. These allowed and excluded regions contradict each other, suggesting that the observed nonlinearity may be induced by some background effects within the SM. This calls for further investigation. A recent observation of nonlinearity in Ca [272] was based on the combination of the precision spectroscopy of Ca+ and Ca14+. This study attributed the observed nonlinearity to the second order mass shift and nuclear polarization, setting most stringent constraint on the new bosons.
@@
When nonlinearity is observed in the King plot, its source must be carefully analyzed. For this purpose, nonlinearity is decomposed into the following two components: , where is the residual of four data points from a linear fit [370] (Note that in Ref. [368] is equivalent to ). On the plane, a single source of nonlinearity with varying strength forms a line passing through the origin. When residuals of King plots for different combination of transitions in Yb are plotted on this plane, the data points approximately align along a single line. This suggests that the observed nonlinearity primarily originates from a single source, which can be explained by nuclear density functional theory calculations. These calculations incorporate several factors, including the surface thickness of nuclear density, relativistic corrections, nuclear deformation, and pairing effects [368]. The isotope shift measurements are so precise that they can select a nuclear model that accurately describes the experimental results. [368, 372, 373].
@@
IX Summary and Outlook
@@
High precision measurements with quantum states in atoms can detect small energy shifts induced by phenomena originating from nuclear and particle physics. For future advancements, several key aspects are important. One is improvement in measurement precision. Even within the same experimental system, increased measurement precision enhances sensitivity to nuclear and particle physics phenomena in most of the topics covered in this paper. In the past, improvements in clock uncertainty contributed to searches for ultralight dark matter and time variation of fundamental constants. Second, the use of highly sensitive quantum sensors can significantly improve the sensitivity. A notable example is searches for variation of the fine structure constant. Switching from conventional ion clocks to HCI clocks or nuclear clocks would enhance the sensitivity by orders of magnitude. In addition, a solid theoretical understanding of the measurements is crucial. Theoretical calculations of atomic or molecular structures determine the sensitivity of a quantum sensor to the quantity of interest. Examples include the sensitivity coefficient for the variation of the fine structure constant. Further improvements in the accuracy of these theoretical calculations can refine the determination of parameters in nuclear and particle physics. Moreover, new theoretical proposals can inspire extensive experimental searches for previously unexplored parameter spaces. Historically, searches for the fifth force using isotope shifts and ultralight dark matter expanded following initial theoretical proposals. With the combination of these aspects, further exploration of nuclear and particle physics using atomic clocks as quantum sensors are expected.
@@ Acknowledgements.
- Cong et al. [2025] L. Cong, W. Ji, P. Fadeev, F. Ficek, M. Jiang, V. V. Flambaum, H. Guan, D. F. Jackson Kimball, M. G. Kozlov, Y. V. Stadnik, et al., “Spin-dependent exotic interactions,” Rev. Mod. Phys. 97, 025005 (2025).
@@
- Marshall et al. [2025] M. C. Marshall, D. A. R. Castillo, W. J. Arthur-Dworschack, A. Aeppli, K. Kim, D. Lee, W. Warfield, J. Hinrichs, N. V. Nardelli, T. M. Fortier, et al., “High-stability single-ion clock with systematic uncertainty,” Phys. Rev. Lett. 135, 033201 (2025).
- Lindvall et al. [2025] T. Lindvall, T. Fordell, K. Hanhijärvi, M. Doležal, J. Rahm, S. Weyers, and A. Wallin, “88Sr+ optical clock with systematic uncertainty and measurement of its absolute frequency with uncertainty,” Phys. Rev. Appl. 24, 044082 (2025).
- Zhang et al. [2025] B. Zhang, Z. Ma, Y. Huang, H. Han, R. Hu, Y. Wang, H. Zhang, L. Tang, T. Shi, H. Guan, et al., “A liquid-nitrogen-cooled Ca+ ion optical clock with a systematic uncertainty of 4.6e-19,” (2025), arXiv:2506.17423 [physics.atom-ph] .
@@
- McGrew et al. [2018] W. F. McGrew, X. Zhang, R. J. Fasano, S. A. Schäffer, K. Beloy, D. Nicolodi, R. C. Brown, N. Hinkley, G. Milani, M. Schioppo, et al., “Atomic clock performance enabling geodesy below the centimetre level,” Nature 564, 87–90 (2018).
@@
- Taichenachev et al. [2006] A. V. Taichenachev, V. I. Yudin, C. W. Oates, C. W. Hoyt, Z. W. Barber, and L. Hollberg, “Magnetic field-induced spectroscopy of forbidden optical transitions with application to lattice-based optical atomic clocks,” Phys. Rev. Lett. 96, 083001 (2006).
@@
- Chen et al. [2025] Z.-A. Chen, H.-R. Zeng, W.-W. Wang, H. Zhang, R.-Q. Lei, J.-Z. Li, C.-Y. Pang, S.-S. Huang, and X. Zhang, “A laser with instability reaching based on a 10-cm-long silicon cavity at sub-5-K temperatures,” Sci. Bull. 70, 3337–3344 (2025).
@@
- Zheng et al. [2022] X. Zheng, J. Dolde, V. Lochab, B. N. Merriman, H. Li, and S. Kolkowitz, “Differential clock comparisons with a multiplexed optical lattice clock,” Nature 602, 425–430 (2022).
@@
- Kim et al. [2025] K. Kim, A. Aeppli, W. Warfield, A. Chu, A. M. Rey, and J. Ye, “Atomic coherence of 2 minutes and instability of at 1 s in a Wannier-Stark lattice clock,” Phys. Rev. Lett. 135, 103601 (2025).
- Ma et al. [2025] S. Ma, J. Dolde, X. Zheng, D. Ganapathy, A. Shtov, J. Chena, A. Stoeltzel, B. J. Christensen, and S. Kolkowitz, “Enhancing optical lattice clock coherence times with erasure conversion,” (2025), arXiv:2505.06437 [physics.atom-ph] .
@@
- Hassan et al. [2025] Y. S. Hassan, K. Beloy, J. L. Siegel, T. Kobayashi, E. Swiler, T. Grogan, R. C. Brown, T. Rojo, T. Bothwell, B. D. Hunt, et al., “Cryogenic optical lattice clock with blackbody radiation stark uncertainty,” Phys. Rev. Lett. 135, 063402 (2025).
@@
- Yang et al. [2025] Y. A. Yang, M. Miklos, Y. M. Tso, S. Kraus, J. Hur, and J. Ye, “Clock precision beyond the standard quantum limit at level,” Phys. Rev. Lett. 135, 193202 (2025).
@@
- Jiang, Arora, and Safronova [2008] D. Jiang, B. Arora, and M. S. Safronova, “Electric quadrupole moments of metastable states of Ca+, Sr+, and Ba+,” Phys. Rev. A 78, 022514 (2008).
@@
- Dietze et al. [2025] K. Dietze, L. Pelzer, L. Krinner, F. Dawel, J. Kramer, N. C. H. Spethmann, T. Kielinski, K. Hammerer, K. Stahl, J. Klose, et al., “Entanglement-enhanced optical ion clock,” (2025), arXiv:2506.11810 [physics.atom-ph] .
- Pezzè et al. [2018] L. Pezzè, A. Smerzi, M. K. Oberthaler, R. Schmied, and P. Treutlein, “Quantum metrology with nonclassical states of atomic ensembles,” Rev. Mod. Phys. 90, 035005 (2018).
@@
- Uzan [2025] J.-P. Uzan, “Fundamental constants: from measurement to the universe, a window on gravitation and cosmology,” Living Rev. Relativ. 28, 6 (2025).
- Webb et al. [1999] J. K. Webb, V. V. Flambaum, C. W. Churchill, M. J. Drinkwater, and J. D. Barrow, “Search for time variation of the fine structure constant,” Phys. Rev. Lett. 82, 884–887 (1999).
@@
- Beeks et al. [2025] K. Beeks, G. A. Kazakov, F. Schaden, I. Morawetz, L. Toscani De Col, T. Riebner, M. Bartokos, T. Sikorsky, T. Schumm, C. Zhang, et al., “Fine-structure constant sensitivity of the Th-229 nuclear clock transition,” Nat. Commun. 16, 9147 (2025).
@@
- Allehabi, Dzuba, and Flambaum [2025] S. O. Allehabi, V. A. Dzuba, and V. V. Flambaum, “Hf12+ ion: Highly charged ion for next-generation atomic clocks and tests of fundamental physics,” (2025), arXiv:2511.00440 [physics.atom-ph] .
@@
- Pfäfflein et al. [2025] P. Pfäfflein, G. Weber, S. Allgeier, Z. Andelkovic, S. Bernitt, A. I. Bondarev, A. Borovik, L. Duval, A. Fleischmann, O. Forstner, et al., “Quantum electrodynamics in strong electromagnetic fields: Substate resolved K transition energies in heliumlike uranium,” Phys. Rev. Lett. 134, 153001 (2025).
- Schippers et al. [2025] S. Schippers, C. Brandau, S. Fuchs, M. Lestinsky, S. X. Wang, C. Y. Zhang, N. R. Badnell, A. Borovik, M. Fogle, V. Hannen, et al., “Testing strong-field qed to second order in the highly correlated atomic system berylliumlike Pb78+ by electron-ion collision spectroscopy,” Phys. Rev. Lett. 135, 113001 (2025).
@@
- Geller [1970] R. Geller, “New high intensity ion source with very low extraction voltage,” Appl. Phys. Lett. 16, 401–404 (1970).
@@
- Cheung et al. [2025] C. Cheung, S. G. Porsev, D. Filin, M. S. Safronova, M. Wehrheim, L. J. Spieß, S. Chen, A. Wilzewski, J. R. C. López-Urrutia, and P. O. Schmidt, “Finding the ultranarrow electric quadrupole transition in Ni12+ ion for an optical clock,” Phys. Rev. Lett. 135, 093002 (2025).
@@
- Rehbehn et al. [2025] N.-H. Rehbehn, L. P. K. Sajith, M. K. Rosner, C. Cheung, S. G. Porsev, M. S. Safronova, S. Worm, D. Budker, T. Pfeifer, J. R. C. López-Urrutia, et al., “Study of the elusive level crossing in highly charged osmium with optical transitions suitable for physics beyond the standard model searches,” (2025), arXiv:2509.06710 [physics.atom-ph] .
- Wilzewski et al. [2025] A. Wilzewski, L. J. Spieß, M. Wehrheim, S. Chen, S. A. King, P. Micke, M. Filzinger, M. R. Steinel, N. Huntemann, E. Benkler, et al., “Nonlinear calcium King plot constrains new bosons and nuclear properties,” Phys. Rev. Lett. 134, 233002 (2025).
@@
- Schaden et al. [2025] F. Schaden, T. Riebner, I. Morawetz, L. T. De Col, G. A. Kazakov, K. Beeks, T. Sikorsky, T. Schumm, K. Zhang, V. Lal, et al., “Laser-induced quenching of the Th-229 nuclear clock isomer in calcium fluoride,” Phys. Rev. Res. 7, L022036 (2025).